Please show all your work and the thought process i.e why you did what you did, also I have to do this by hand with a calculator, so please DON'T use graphs or plot anything. Thank you.

The following data were obtained in a study of an enzyme known to follow Michaelis-Menten kinetics:

V0 (μmol/min)

—————————————

217

325

433

488

647

—————————————

Substrate added (mmol/L):

0.8

2

4

6

1,000

The Km for this enzyme is approximately:

A) 1 mM

B) 2 mM

C) 4 mM

D) 6 mM

E) 1,000 mM

Thank you for all your help. The correct answer is 2mM
Answer #1

V0 = 217

S = 0.8

V = Vmax (S)/Km*(S)

V0 =Vmax/2

Vmax = V0 * 2

         = 217 *2

         = 434

Km = Vmax *(S)/(S) *V0

        = 434*0.8/0.8*217

      = 347.2/173.6 = 2mM